40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 27 18:03:47 2014, ericq, Update, Green Locking, uPDH Box Checkup uPDH_X.JPGuPDH_Y.JPGupdhTFs.pdfupdhNoises.pdfuPDH_Aug27.zip
    Reply  Thu Aug 28 17:34:20 2014, ericq, Update, Green Locking, uPDH Box Checkup TEK00006.PNG
       Reply  Fri Aug 29 17:28:07 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Fri Aug 29 23:53:35 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Tue Sep 2 20:02:06 2014, ericq, Update, Green Locking, uPDH Box Checkup 
       Reply  Wed Sep 3 00:17:22 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Wed Sep 3 20:38:40 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Fri Sep 5 04:07:44 2014, ericq, Update, Green Locking, X end uPDH Box Replaced 
                Reply  Fri Sep 5 16:13:00 2014, ericq, Update, Green Locking, new X Green PDH measurements 6x
                   Reply  Sat Sep 6 01:48:54 2014, rana, Update, Green Locking, new X Green PDH measurements 
                      Reply  Sat Sep 6 14:49:12 2014, ericq, Update, Green Locking, new X Green PDH measurements xGComparison.pdfxGspectra.pdf
                         Reply  Sun Sep 7 17:06:38 2014, rana, Update, Green Locking, new X Green PDH measurements 
                         Reply  Mon Sep 8 14:14:13 2014, Jenne, Update, Green Locking, new Y Green PDH measurements 6x
Message ID: 10471     Entry time: Mon Sep 8 14:14:13 2014     In reply to: 10464
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: new Y Green PDH measurements 

These are plots and notes from last week's PDH adventures. 


For the PDH servo box re-design, we wanted to think a little bit about what we actually wanted out of the box.

* We want the zero of the main transfer function to be at the same frequency as the cavity pole for green, which is about 18kHz.

* We want the boost to suppress noise at a few hundred Hz.  We don't need super-duper low-frequency boost, nor do we want it.  We'd like to leave the boost on all the time.

* Wanted to get rid of 10dB attenuator on PD input, so needed to lower the overall gain.

* We acknowledge that the gain of the raw error signal times the PZT response is very high, so no matter what, we will have to have a low-gain servo, even perhaps have the servo shape be less than unity gain.

---> We reduced the gain of the first amplification stage from a gain of 20 to a gain of 3.

---> Made the boost stage have a DC gain of 1.  Pole at 75 Hz and Zero at 1.6kHz to give suppression at a few hundred Hz.  Boost is *not* a pure integrator, so that we can leave it on.  (If we required triggering anyway, we would have made it a pure integrator).

---> In transfer function stage, put zero at 17.7kHz to match cavity pole.  Pole of servo was going to be at 20 Hz, but we wanted a little more gain, so we lowered it to 2 Hz.

Here is the final measured servo box transfer function for the Yend box (with an arbitrary gain knob setting):

PDHboxTF.png


Once installed, I set the gain knob for the Yend at 4.0, which gave an overall UGF of about 10kHz.  Then I measured the loop:

LoopGain.png

I also measured the error point and the control point, and compared them to Q's measurements in elog 10430.

ErrSpecCompare.png

ControlSpectrumCompare.png


In order to see what we might expect for a contribution to ALS noise, I looked at the error point spectra and lowpassed it with a pole at 200Hz.  I do this because the PDH error is like sensor noise for the ALS, but the ALS UGF is around 200 Hz, so noise at frequencies higher than that will be suppressed like 1/f.  So, I lowpass the error signal, then look at the RMS, and see that we should be pretty happy with our result. I include also the Xend error spectrum, as measured and reported by Q in elog 10460.

XvsY_ALScontribution.png

Attachment 6: Yend_4Sept2014.zip  2.515 MB  Uploaded Mon Sep 8 15:26:43 2014
ELOG V3.1.3-